# ART-DECO: Arbitrary Text Guidance for 3D Detailizer Construction

### v1

@@ Appendix B Network architecture
For the density upsampling network, we use 5 layers of 3D convolution to extract the features of the input coarse voxel grid, followed by 2 layers of transposed 3D convolution for upsampling. Each upsampling layer doubles the input resolution. For the albedo upsampling network, we use the same network architecture as the density upsampling network, except that the final upsampling layer outputs three channels to represent RGB values.
We use a learning rate of $10^{-4}$, a batch size of $1$ and Adam optimizer for all experiments.
@@ Appendix E ShaDDR implementation
Since our method generates the detailed shape is in radiance field representation, for a fair comparison, we modify one of the ablation settings in ShaDDR to also produce a radiance field, ensuring a fair comparison. More specifically, we modify the supervision of ShaDDR’s texture branch from a 2D discriminator to a 3D discriminator, which can directly discriminate the output albedo field against the albedo field of the style shape. We provide the detailed shapes generated by our method, which are used as styles for training ShaDDR in Figure 17.
Appendix F Additional quantitative comparisons
Table 3 to 10 show the quantitative comparisons of randomly selected text prompts in our experiments.
Appendix G Additional quantitative ablations
Table 11 to 18 show the quantitative ablation of using regularization loss with different $\lambda_{reg}$ on randomly selected text prompts.
Appendix H Additional qualitative results

### v3

@@ Appendix B Network architecture
We took inspiration from ShaDDR (Chen et al., 2023b), where the input voxel grid is dilated by one voxel and two branches of networks are used to generate geometry and texture, respectively. For the density upsampling network, we use 5 layers of 3D convolution to extract the features of the input coarse voxel grid, followed by 2 layers of transposed 3D convolution for upsampling. Each upsampling layer doubles the input resolution. For the albedo upsampling network, we use the same network architecture as the density upsampling network, except that the final upsampling layer outputs three channels to represent RGB values.
We use a learning rate of $10^{-4}$, a batch size of $1$, and the Adam optimizer for all experiments.
@@ Appendix E ShaDDR implementation
We use the original implementation of ShaDDR and its default settings: an input resolution of $k=32$ and an output resolution of $K=256$. We provide the detailed shapes generated by our method as styles for training ShaDDR in Figure 19.
Appendix F Geometry-only visualizations
Figure 15 shows the geometry-only visualizations corresponding to Figure 12 in the main paper.
Appendix G Additional qualitative comparisons
Figure 16 shows additional comparisons with Trellis and Sherpa3D. Trellis tends to preserve only the input coarse structure and fails to add geometric or texture details, resulting in outputs that do not match the style of the input text prompt. Sherpa3D tends to disregard the coarse initialization and instead optimizes the geometry to match the priors of its 2D diffusion model. For example, given a coarse structure of an animal with two heads and six legs, Sherpa3D simplifies it and generates one head and four legs. The structures of chairs, tables, and buildings are also overly simplified.
Appendix H User Study
For the user study, we used the same examples that were used to compute the quantitative results. We designed 40 questions and collected responses from 42 participants. For each question, we show the participants an input coarse shape for structural control, a text prompt describing the style, and several 3D shapes generated by our method and three baselines. We ask the participants to select the BEST generated 3D shape based on several aspects:
  • Style adherence - How well the generated 3D shape follows the style described by the text.
  • Quality of the generated details - As judged by the fine-grained details of the generated 3D shape.
  • Structure preservation - How well the generated shape follows the structure of the input coarse shape.
  • Overall preference - As judged by the overall impression of the generation, all factors considered.
We report the average percentage of participants selecting each method as producing the best results. As shown in Table 3, 83.45% of the participants selected the shapes generated by our method as the most preferred result, highlighting its advantage over competing approaches.
Table 3. User study results. The average percentage of participants selecting each method as producing the best results.
Ours CLAY Coin3D ShaDDR
Avg. preference 83.45% 15.77% 0.06% 0.72%
Appendix I Additional quantitative comparisons
Table 4 to $1$ show the quantitative comparisons of randomly selected text prompts in our experiments.
Appendix J Additional quantitative ablations
Table 12 to 19 show the quantitative ablation of using regularization loss with different $\lambda_{reg}$ on randomly selected text prompts.
Appendix K Additional qualitative results
@@ Appendix I Additional quantitative comparisons
Refer to caption“a cute smooth-coatedLabrador retriever”“a wild coyote”“a short-hairedtabby cat”“a wild grizzly bear”“a largeBengal tiger”“a robot dog”“a Gothic church withspires, detailed carvingsand flying buttresses”“an Italian basilicawith domed roof“a medieval castle withtowers, battlementsand stone walls”“a large stone castlewith round towersand pointed arches”“a castle with cone-shaped tower roofsand brick walls”“a Byzantine cathedralcathedral with talldomes, mosaics”“a folding tablewith plastic surfaceand metal frame”“an industrial tablewith a metal frameand a wood top”“a mid-century moderntable with clean linesand angled legs”“a luxury marbletable with astone surface”“a stone tablewith a heavy baseand rough texture”“a table with metalframe and a simplewood surface”
Figure 15. Geometry-only visualization of text-guided detailization with input coarse voxels control. We show the input coarse voxels on the left and the text prompts on the top.
Ours
Sherpa3D
Trellis
Ours
Sherpa3D
Trellis
“an industrial table with ametal frame and wood top”“a Queen Anne table withsmooth and muted finish”“a cute bulldog”“a cute Shiba Inu”“a cute golden retriever”“an office chair withwheels and thick padding”“a rustic wooden chairwith rough texture look”“an Italian basilicawith domed roof”“a castle with cone-shapedtower roofs and brick walls”
Figure 16. Additional comparison results with TRELLIS and Sherpa3D.
@@
Refer to caption(a) Rendering of the input coarse proxy(b) Single-view image(c) Multi-view images(d) Rendering of NeuS reconstruction from (c)